PLEASE I NEED IT ASAP! A man is 27 years older than his son and in 10 years, he will be twice as old as his son. How old is each now?
Answer by ankor@dixie-net.com(22740)

A man is 27 years older than his son and in 10 years, he will be twice as old as his son. How old is each now?
:
Let m = man's present age
Let s = son's present age
:
Write an equation for each statement:
:
"A man is 27 years older than his son"
m = s + 27
:
"in 10 years, he will be twice as old as his son."
m + 10 = 2(s+10)
m + 10 = 2s + 20
m = 2s + 20 - 10
m = 2s + 10
:
From the 1st statement, replace m with (s+27)
s + 27 = 2s + 10
27 - 10 = 2s - s
17 = s
:
Son is 17 years old, I'll let you find the man's age, check your solutions in the 2nd statement
:
Hey! Try not to wait til the last minute! C